Draw graphs of the following functions.

(a)

f: R—>R defined by f(x) = x

(b)

f: R—>R defined by f(x) = |x|

(c)

f: R—>R defined by f(x) = x + 1

(d)

f: R—>R defined by f(x) = – 3x+4

(e)

f: R—>R defined by f(x) = Floor(x)

(f)

f: R—>R defined by f(x) = Ceiling(x)

(g)

f: R—>R defined by f(x) = x2

(h)

f: R—>R defined by f(x) = x3


1
Expert's answer
2021-07-20T04:24:21-0400

(a) "f:R\rightarrow R" defined by "f(x) = x"





(b) "f:R\rightarrow R" defined by "f(x) = |x|"





(c) "f:R\rightarrow R" defined by "f(x) = x+1"





(d) "f:R\rightarrow R" defined by "f(x) = -3x+4"




(e) "f:R\rightarrow R" defined by "f(x) = Floor(x)"


The floor function is the function that takes as input a real number "x," and gives as output the greatest integer less than or equal to "x," denoted "Floor(x)" or "\lfloor x \rfloor."




(f) "f:R\rightarrow R" defined by "f(x) = Ceiling(x)"


 The ceiling function is the function that takes as input a real number "x," and gives as output the least integer greater than or equal to "x," denoted "Ceiling(x)" or "\lceil x\rceil."




(g) "f:R\rightarrow R" defined by "f(x) = x^2"





(h) "f:R\rightarrow R" defined by "f(x) = x^3"